# FissionFusion: Fast Geometric Generation and Hierarchical Souping for Medical Image Analysis

### v1

@@ 2 Methodology
[…]
@@ 3 Experiments
Implementation Details. Our research investigates two model architectures: DeiT-S (Transformer) and ResNet50 (CNN), both pre-trained on ImageNet. Before full fine-tuning, we conduct a 10-epoch warmup where the whole network except the last layer is frozen ($\theta_{\text{0}}$). We then unfreeze the entire model and perform a full fine-tuning known as the LP-FT approach [15]. We use AdamW optimizer with a cosine annealing scheduler for 50 epochs, batch size of 128, and an image resolution of 224$\times$224.
@@ 4 Results and Discussion
Figure 3: OOD analysis for different architectures on various datasets (a) CIFAR10 v/s CIFAR10.1 - ResNet50 (b) CheXpert v/s MIMIC - ResNet50 (c) APTOS v/s (EyePacs, Messidor, Messidorv2) - ResNet50 (d) CIFAR10 v/s CIFAR10.1 - DeiT-S (e) CheXpert v/s MIMIC - DeiT-S (f) APTOS v/s (EyePacs, Messidor, Messidorv2) - DeiT-S. We do not plot the results of Uniform Soups as it performs poorly.
As seen in Table 1, we observe that our method achieves better results both in the natural as well as medical imaging domains. In CIFAR datasets, we observe that GoU and GoG achieve better results compared to model soups, though it is not significant due to the smooth convex error surface. However, when testing on medical datasets, we observe around 6% improvement in Recall (GoG) for the HAM10000 dataset (ResNet50) as well as around 6% gain (GoG) in AUC(DeiT-S) for the CheXpert dataset. Both the GoU as well as GoG approaches achieve similar results except in the case of CheXpert (DeiT-S), where GoU significantly lags behind GoG. The GoG is always better than the Greedy Soup approach in almost all cases and is achieved at a significantly lower computational cost, as we do not have to perform the full grid search. Greedy soups does not perform as expected when fine-tuning on medical datasets as the averaging starts from the best model, which can be an issue due to uneven error surfaces. For example, for the DeiT-S on the CheXpert dataset, we observe that the best model has a high validation score, but a low test score indicating its not generalizable as it might have been stuck in a local minima. The FGG process overcomes this issue through its ability to escape numerous local minima, while the acquired models cluster around the same area in the error surface, facilitating smoother averaging. The local averaging in HS approach allows smoother averaging at local surfaces and greater diversity at the top level. We further conduct an ablation study in Table 1 in Section A1 in the Appendix, exploring greedy souping on the FGG models, and HS on the grid-search generated models.

### v2

@@ 2 Methodology
In FGG, we start with $\theta_{\text{0}}$ as an initialization and fully train the model using $n$ different learning rates, resulting in a set of parameters $\Theta=\{\theta_{1},\dots,\theta_{n}\}$. After obtaining $\Theta$, we implement the Fission process. We initialize the weights $\theta_{t}\in\Theta$, and carry out a second training process for a set of iterations $\textit{I}=\{1,\dots,k\}$ to generate a set of parameters $\hat{\Theta}=\{\hat{\theta}^{1},\dots,\hat{\theta}^{m}\}$. During this, we employ a learning rate scheduler to cyclically change the learning rate every cycle $c$, where $c$ is defined as a set of number of iterations and is an even number. This encourages $\theta_{t}$ to diverge from its initialization, while maintaining validation accuracy. We use a function $\alpha(i)$ to control the learning rate at iteration $i\in\textit{I}$ as follows:
@@ 3 Experiments
Implementation Details. Our research investigates two model architectures: DeiT-B (Transformer) and ResNet50 (CNN), both pre-trained on ImageNet. Before full fine-tuning, we conduct a 10-epoch warmup where the whole network except the last layer is frozen ($\theta_{\text{0}}$). We then unfreeze the entire model and perform a full fine-tuning known as the LP-FT approach [15]. We use AdamW optimizer with a cosine annealing scheduler for 50 epochs, batch size of 128, and an image resolution of 224$\times$224.
@@ 4 Results and Discussion
Figure 3: OOD analysis for different architectures on various datasets (a) CIFAR10 v/s CIFAR10.1 - ResNet50 (b) CheXpert v/s MIMIC - ResNet50 (c) APTOS v/s (EyePacs, Messidor, Messidorv2) - ResNet50 (d) CIFAR10 v/s CIFAR10.1 - DeiT-B (e) CheXpert v/s MIMIC - DeiT-B (f) APTOS v/s (EyePacs, Messidor, Messidorv2) - DeiT-B. We do not plot the results of Uniform Soups as it performs poorly.
As seen in Table 1, we observe that our method achieves better results both in the natural as well as medical imaging domains. In CIFAR datasets, we observe that GoU and GoG achieve better results compared to model soups, though it is not significant due to the smooth convex error surface. However, when testing on medical datasets, we observe around 6% improvement in Recall (GoG) for the HAM10000 dataset (ResNet50) as well as around 6% gain (GoG) in AUC(DeiT-B) for the CheXpert dataset. Both the GoU as well as GoG approaches achieve similar results except in the case of CheXpert (DeiT-B), where GoU significantly lags behind GoG. The GoG is always better than the Greedy Soup approach in almost all cases and is achieved at a significantly lower computational cost, as we do not have to perform the full grid search. Greedy soups does not perform as expected when fine-tuning on medical datasets as the averaging starts from the best model, which can be an issue due to uneven error surfaces. For example, for the DeiT-B on the CheXpert dataset, we observe that the best model has a high validation score, but a low test score indicating its not generalizable as it might have been stuck in a local minima. The FGG process overcomes this issue through its ability to escape numerous local minima, while the acquired models cluster around the same area in the error surface, facilitating smoother averaging. The local averaging in HS approach allows smoother averaging at local surfaces and greater diversity at the top level. We further conduct an ablation study in Table 1 in Section A1 in the Appendix, exploring greedy souping on the FGG models, and HS on the grid-search generated models.